OpenStudy (anonymous):

solve for x 3(X-5)-4X = 6(X-10) -(X+ 5)

OpenStudy (anonymous):

20/3

OpenStudy (anonymous):

3(X-5)-4X = 6(X-10) -(X+ 5) 3X-15-4x=6X-60-X-5 -7X-15=5X-65 -12X=-50 X=50/12

OpenStudy (anonymous):

sorry 3X-15-4x=6X-60-X-5 -x-15=5X-65 -6X=50 X=-50/6

OpenStudy (anonymous):

its +5 not -5

OpenStudy (anonymous):

actually -(X+5)-- when distributed --- becomes --- -X-5 so 3(X-5)-4X = 6(X-10) -(X+ 5) 3X-15-4x=6X-60-X-5 -x-15=5X-65 -X-5X= -65+15 -6X=-50 X=50/6

OpenStudy (anonymous):

since its in the bracket the negatives sign gets distributed

OpenStudy (anonymous):

The minus sign doesnt represent -1 in this case, it is the subtraction of the sum of X + 5, isnt it?

OpenStudy (anonymous):

distributive property says if there is a for example 10M-(5X+5) it will be 10M -5X-5